OpenStudy (deadshot):

What is the period of the function f(x)=cos4x?

ganeshie8 (ganeshie8):

period of \(\cos(x) = 2\pi\) so, to find period of \(\cos(4x)\) simply equate \(4x\) to \(2pi\) \(4x = 2\pi\) solve \(x\)

OpenStudy (deadshot):

So divide both sides by 4 to get \[x=\pi\]

ganeshie8 (ganeshie8):

\(4x = 2\pi\) \(x = \frac{2\pi}{4}\)

OpenStudy (deadshot):

Can I hire you through paypal as a Tutor? I can't afford to spend the next 5 days on this 25 question exam.

OpenStudy (deadshot):

so it would be \[x=\frac{ 1 }{ 2 }\pi\]?

ganeshie8 (ganeshie8):

lol that wont be necessary, just tag me in your questions, il try to help :)

ganeshie8 (ganeshie8):

Yup !

ganeshie8 (ganeshie8):

period of \(\cos(4x)\) is \(\frac{\pi}{2}\)

OpenStudy (deadshot):

Thanks!

ganeshie8 (ganeshie8):

np :)